OpenStudy (anonymous):

use log 4= 0.062 and log 12= 1.079 to evaluate the logarithm. 12.) log 1/3

OpenStudy (anonymous):

\[\log(\frac{1}{3}) = \log(\frac{4}{12}) = \log(4) - \log(12) = 0.062 - 1.079 = -1.017\]

OpenStudy (anonymous):

thank you :)